OpenStudy (anonymous):

what is the simplified form of -9m^-2n^5*2m^-3m^-6

OpenStudy (anonymous):

\[-9m ^{-2}n ^{5}*2m ^{-3}n ^{-6}\]

OpenStudy (anonymous):

I think its \[-18m ^{6}n ^{30}\] but im not sure

OpenStudy (jordanloveangel):

-9m(-2)n5*2m(-3)m(-6) Final result : ( -32•2m(-11)n5) Reformatting the input : Changes made to your input should not affect the solution: (1): "^-6" was replaced by "^(-6)". 2 more similar replacement(s) Step by step solution : Step 1 : Multiply (32•2m(-2)n5) by m(-3) Multiplying exponential expressions : 1.1 m(-2) multiplied by m(-3) = m((-2) + (-3)) = m(-5) = 1/m5 Equation at the end of step 1 : 0 - ((32•2m(-5)n5) • m(-6)) Step 2 : Multiply (32•2m(-5)n5) by m(-6) Multiplying exponential expressions : 2.1 m(-5) multiplied by m(-6) = m((-5) + (-6)) = m(-11) = 1/m11 Equation at the end of step 2 : 0 - (32•2m(-11)n5) Step 3 : Simplify 0 - (32•2m(-11)n5) Final result : ( -32•2m(-11)n5)

OpenStudy (anonymous):

thank you

OpenStudy (jordanloveangel):

did it help doe?